OpenStudy (anonymous):

write an expression as the cosine of an angle, knowing that the expression is the right side of the formula for cos(a-b) with particular values for (a-b) cos(165)cos(45)+sin(165)sin(45)

OpenStudy (goformit100):

A warm Welcome to OpenStudy. I can help guide you through this useful site. You can ask your questions to me or you can message me. Please use the chat for off topic questions. Remember to give the person who helped you a medal by clicking on "Best Answer". We follow a code of conduct, ( http://openstudy.com/code-of-conduct ). Please take a moment to read it.

OpenStudy (gorv):

cos (a-b)=cosa*cosb+sina*sinb cos(165)cos(45)+sin(165)sin(45)=cos(165-45)=cos120=-1/2